Calculate the yearly interest rate if an investment is paid 1.75% interest every two months 17 2.18 Calculate the interest rate per interest period if the yearly interest rate is 13% and the number of interest periods per year is three. 2.19 Calculate the number of interest periods per year if the yearly interest rate is 15% and the interest rate per interest period is 2.5% 2.20 Calculate the yearly interest rate if there are 12 interest periods per year and the interest rate per period is 0.8% 2.21 Calculate the effective interest rate if the interest rate per compounding period is 1.5% and the number of compounding periods per year is 12. 2.22 Calculate the effective interest rate if the nominal interest rate is 16% and the interest is compounded every two months. 2.23 1f a bank pays 9%interest per year and the compounding period is weekly, what is the effec- tive interest rate? 2.24 An engineer locates a bank paying 2% interest per year compounded continuously. If the engi- neer invests his money in this bank, what is the continuously compounded interest rate? 2.25 An engineering firm borrows funds from a bank charging 9% interest compounded continu- ously. What is the continuously compounded interest rate? If a credit union pays 1% yearly compounded monthly, what is the effective interest rate? 2.26 2.27 If an engineer is earning 1.75% interest per year compounded daily, what is the effective 2.28 Calculate the nominal interest rate if the period interest rate is 0.5% and there are four com- 2.29 A firm pays year-end dividends to its employees of 1.5% but the funds in the account com- 2.30 What is the continuously compounded interest rate if an investor is able to invest her funds interest rate? pounding periods per year. pound every two months. What is the effective interest rate? at a nominal interest rate of 1.25%? I need 17-30!!
Answer #1

(2.17)

Yearly nominal interest rate = 1.75% x (12/2) = 10.5%

Yearly Effective interest rate (EAR) = [1 + r]N - 1, where N: Number of compounding periods

r = 1.75% = 0.0175, N = 12/2 = 6

EAR = [1 + (0.0175]6 - 1 = (1.0175)6 - 1 = 1.1097 - 1 = 0.1097 = 10.97%

(2.18)

Nominal periodic interest rate = 13% / 3 = 4.33%

(2.19)

Number of interest periods = Annual interest rate / Periodic interest rate = 15% / 2.5% = 6

(2.20)

Yearly nominal interest rate = 0.8% x 12 = 9.6%

Yearly Effective interest rate (EAR) = [1 + r]N - 1, where N: Number of compounding periods

r = 0.8% = 0.008, N = 12

EAR = [1 + (0.008]12 - 1 = (1.008)12 - 1 = 1.1003 - 1 = 0.1003 = 10.03%

NOTE: As per Answering Policy, 1st 4 questions are answered.